OpenStudy (anonymous):

Solve: –2x – 15 = 6x + 9

OpenStudy (anonymous):

-2x-15=6x+9 -8x=24 -x=3 x=-3

OpenStudy (anonymous):

-2x-15=6x+9 -8x=24 -x=3 x=-3